OpenStudy (jellybot23):

HELP PLEASEEEE write the expression 3log(x-2)+log(x+2)-log(x^2-4) as single logarithm

OpenStudy (michele_laino):

Here, we can apply these properties of logarithms: \[\huge \begin{gathered} k\log \left( x \right) = \log \left( {{x^k}} \right) \hfill \\ \hfill \\ \log \left( x \right) + \log \left( y \right) = \log \left( {xy} \right) \hfill \\ \hfill \\ \log \left( x \right) - \log \left( y \right) = \log \left( {\frac{x}{y}} \right) \hfill \\ \end{gathered} \] being \(k\) any real number, and \(x>0,\;y>0\)

OpenStudy (jellybot23):

Okay, yes I know that rules.. I think I know the answer but am not sure. Should I multiply 3log(x-2)+log(x+2)? Or should I turn 3log(x-2) into log(x-2^3) first?

OpenStudy (jellybot23):

@Michele_Laino

OpenStudy (jellybot23):

come back please

OpenStudy (michele_laino):

first step: I apply this property: \[\huge k\log \left( x \right) = \log \left( {{x^k}} \right)\] so, I can write: \[\huge 3\log \left( {x - 2} \right) = \log \left\{ {{{\left( {x - 2} \right)}^3}} \right\}\]

OpenStudy (jellybot23):

okay that's what i did too :)

OpenStudy (michele_laino):

now, we have to apply the subsequent property: \[\huge \log \left( a \right) + \log \left( b \right) = \log \left( {ab} \right)\] to this expression: \[\huge \log \left\{ {{{\left( {x - 2} \right)}^3}} \right\}\log \left( {x + 2} \right) = ...?\] please continue

OpenStudy (michele_laino):

oops.. to this expression: \[\huge \log \left\{ {{{\left( {x - 2} \right)}^3}} \right\} + \log \left( {x + 2} \right) = ...?\]

OpenStudy (jellybot23):

That would turn into log(x^2-4)^3??

OpenStudy (michele_laino):

no, we have this: \[\huge a = {\left( {x - 2} \right)^3},\;b = \left( {x + 2} \right)\] so, we have: \[\huge \begin{gathered} \log \left\{ {{{\left( {x - 2} \right)}^3}} \right\} + \log \left( {x + 2} \right) = \hfill \\ = \log \left\{ {{{\left( {x - 2} \right)}^3}\left( {x + 2} \right)} \right\} \hfill \\ \end{gathered} \]

OpenStudy (jellybot23):

oh okay, so you wouldn't multiply the bases together?

OpenStudy (michele_laino):

please wait, since we have to complete the starting expression. Now I apply this property: \[\huge \log \left( a \right) - \log \left( b \right) = \log \left( {\frac{a}{b}} \right)\] to this expression: \[\Large \log \left\{ {{{\left( {x - 2} \right)}^3}\left( {x + 2} \right)} \right\} - \log \left( {{x^2} - 4} \right)\] where: \[\huge a = {\left( {x - 2} \right)^3}\left( {x + 2} \right),\;b = \left( {{x^2} - 4} \right)\] what do you get?

OpenStudy (jellybot23):

I got log[(x-2)^3(x+2)] / log(x^2-4)

OpenStudy (michele_laino):

I got a different result: \[\Large \begin{gathered} \log \left\{ {{{\left( {x - 2} \right)}^3}\left( {x + 2} \right)} \right\} - \log \left( {{x^2} - 4} \right) = \hfill \\ \hfill \\ = \log \left\{ {\frac{{{{\left( {x - 2} \right)}^3}\left( {x + 2} \right)}}{{\left( {{x^2} - 4} \right)}}} \right\} \hfill \\ \end{gathered} \] am I right?

OpenStudy (jellybot23):

Oh yes, because it is a single logarithm... I have never understood logs, so I'm sorry for all the confusion hehe

OpenStudy (michele_laino):

ok! :) Now please keep in mind that we have the subsequent identity: \((x^2-4)=(x-2)(x+2)\), so we can rewrite the last step as below: \[\Large \begin{gathered} \log \left\{ {{{\left( {x - 2} \right)}^3}\left( {x + 2} \right)} \right\} - \log \left( {{x^2} - 4} \right) = \hfill \\ \hfill \\ = \log \left\{ {\frac{{{{\left( {x - 2} \right)}^3}\left( {x + 2} \right)}}{{\left( {{x^2} - 4} \right)}}} \right\} = \hfill \\ \hfill \\ = \log \left\{ {\frac{{{{\left( {x - 2} \right)}^3}\left( {x + 2} \right)}}{{\left( {x - 2} \right)\left( {x + 2} \right)}}} \right\} \hfill \\ \end{gathered} \] now, we can simplify, the expression inside the parenthese, what do you get?

OpenStudy (michele_laino):

parentheses*

OpenStudy (michele_laino):

when you have written the answer, please tag

OpenStudy (jellybot23):

Would it just be log{(x-2)^3 / (x-2)}

OpenStudy (jellybot23):

@Michele_Laino

OpenStudy (michele_laino):

such expression, can be simplified further: so, we have the subsequent steps: \[\Large \begin{gathered} \log \left\{ {{{\left( {x - 2} \right)}^3}\left( {x + 2} \right)} \right\} - \log \left( {{x^2} - 4} \right) = \hfill \\ \hfill \\ = \log \left\{ {\frac{{{{\left( {x - 2} \right)}^3}\left( {x + 2} \right)}}{{\left( {{x^2} - 4} \right)}}} \right\} = \hfill \\ \hfill \\ = \log \left\{ {\frac{{{{\left( {x - 2} \right)}^3}\left( {x + 2} \right)}}{{\left( {x - 2} \right)\left( {x + 2} \right)}}} \right\} = \hfill \\ \hfill \\ = \log \left\{ {{{\left( {x - 2} \right)}^2}} \right\} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

hint: